Present value: Cheryl wants to have 4,600$ in spending money to take on a trip to Disney world in 3 years. How much must she deposit now in a savings account that pays 4% to have the money she needs in three years?
To have 4600$ in 3 years Cheryl would need to deposit ? $_______

Answer by Boreal(15235)

4600=Po{1+0.04)^3
ln of both sides
8.4338=3ln(1+0.04)+ln (Po)
add logs when the numbers are being multiplied
8.4338=0.11766+ln(Po)
8.3161=ln Po.
Raise to the e power and now round (I rounded here, but keep it unrounded on calculator)
$4089.38
4089.38(1.04)^3=
(1.04^3=1.1248
4089.38*1.1248=$4600.